The demand for tickets to an Ethiopian Camparada film is given by D(p)= 200,000-


10,000p, where p is the price of tickets.If the price of tickets is 12 birr, calculate price


elasticity of demand for tickets and draw the demand curve?

1
Expert's answer
2021-11-07T19:40:21-0500

Elasticity of demand

"=\frac{dq}{dP}\times\frac{P}{q}"

"=\frac{d(200,000-10,000P)}{dP}\times \frac{P}{(200,000-10,000P)}"

"=\frac{-P}{20-P}"

Substituting the value of P:

"=\frac{-12}{20-12}"

"=1.5"